OpenStudy (anonymous):

I need help factoring x squared minus 12x + 3

jimthompson5910 (jim_thompson5910):

You need to find two numbers that multiply to 3 and add to 12

jimthompson5910 (jim_thompson5910):

oops meant -12

OpenStudy (anonymous):

-1, 0 ??

jimthompson5910 (jim_thompson5910):

No

OpenStudy (anonymous):

-3, -6?

jimthompson5910 (jim_thompson5910):

No, (-3)*(-6) = 18, but we want 3

OpenStudy (anonymous):

1,-3

jimthompson5910 (jim_thompson5910):

closer, but 1*(-3) = -3, and we want positive 3

OpenStudy (anonymous):

1 and 3?

jimthompson5910 (jim_thompson5910):

better, but there's a problem, does 1 and 3 add to -12?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):

so it's not 1 and 3. The same can be said about -1 and -3

OpenStudy (anonymous):

-4 and -3?

jimthompson5910 (jim_thompson5910):

no because -4 times -3 is NOT 3

jimthompson5910 (jim_thompson5910):

we've basically run out of options because the only factors of 3 are 1 and 3 so because we've exhausted everything and still haven't found the two numbers, this means that x^2-12x+3 can't be factored

OpenStudy (anonymous):

hmmm I have to solve this equation (x-4) squared equals 2x squared minus 4x+19

OpenStudy (anonymous):

correct, kind of like 0 lol

OpenStudy (anonymous):

and solve for x

jimthompson5910 (jim_thompson5910):

what do you get when you expand out (x-4)^2

OpenStudy (anonymous):

x squared + 16

OpenStudy (anonymous):

is that right?

jimthompson5910 (jim_thompson5910):

close, but there should be a middle x term

OpenStudy (anonymous):

x squared -4x+16?

OpenStudy (anonymous):

I get stuck on the middle x term when I'm expanding

jimthompson5910 (jim_thompson5910):

Complete steps and answer for expanding (x-4)^2: (x-4)^2 = (x-4)(x-4) (x-4)^2 = x(x-4)-4(x-4) (x-4)^2 = x*x+x*(-4)-4*x-4*(-4) (x-4)^2 = x^2 - 4x - 4x + 16 (x-4)^2 = x^2 - 8x + 16 So (x-4)^2 expands out to x^2 - 8x + 16

jimthompson5910 (jim_thompson5910):

So (x-4)^2 = 2x^2 - 4x+1 turns into x^2 - 8x + 16 = 2x^2 - 4x + 1

jimthompson5910 (jim_thompson5910):

Let me know if that helps or not.

OpenStudy (anonymous):

ya that helped. so now the equation would be 2x^2 - 4x - 8x+16+19?

jimthompson5910 (jim_thompson5910):

well if it were an equation, there'd be an equal sign somewhere...not sure where you placed yours in that last expression.

OpenStudy (anonymous):

oh it's 0 equals that equation

OpenStudy (anonymous):

so now I just need to factor that equation, so I can solve for x

jimthompson5910 (jim_thompson5910):

You should have x^2 - 8x + 16 = 2x^2 - 4x+19 0 = 2x^2 - 4x+19 - x^2 + 8x - 16 0 = x^2 + 4x + 3

OpenStudy (anonymous):

so then x is -1 and -3

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

thanks so much for your help

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

could you check if I did this right? I have (v+5)^2 which needs to be expanded. I got v^2+50x+25

OpenStudy (anonymous):

and if I didn't could you explain what I did wrong?

jimthompson5910 (jim_thompson5910):

you have the middle term incorrect

jimthompson5910 (jim_thompson5910):

try thinking of (v+5)^2 as (v+5)(v+5)

jimthompson5910 (jim_thompson5910):

then expanding that to get v(v+5)+5(v+5)

OpenStudy (anonymous):

would the middle term be 10v?

jimthompson5910 (jim_thompson5910):

bingo

OpenStudy (anonymous):

so now I have to factor out 0 equals 2v^2 + 14v+29+10v+25

jimthompson5910 (jim_thompson5910):

so combine like terms on the right side, what do you get when you do that?

OpenStudy (anonymous):

2v^2+24v+54

OpenStudy (anonymous):

the gcf is 6 I believe

jimthompson5910 (jim_thompson5910):

no, 6 does not go into 2v^2

OpenStudy (anonymous):

then is it 2?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

factor out 2 to get???

OpenStudy (anonymous):

so I have (v+4) (2v+)

jimthompson5910 (jim_thompson5910):

if you factor out a 2, you'll go from 2v^2+24v+54 to 2(v^2+12v+27)

jimthompson5910 (jim_thompson5910):

so you can focus on factoring v^2+12v+27

OpenStudy (anonymous):

(v+9)(v+3)

jimthompson5910 (jim_thompson5910):

good, so the complete factorization is 2(v+9)(v+3)

OpenStudy (anonymous):

that didn't work...when I entered the answers, it came out wrong

jimthompson5910 (jim_thompson5910):

what was the original problem

OpenStudy (anonymous):

(v+5)^2=2v^2+14v+29

jimthompson5910 (jim_thompson5910):

oh i see where you went wrong, you somehow added 10v to 14v when you should have subtracted 10v from 14v

OpenStudy (anonymous):

ohhh

jimthompson5910 (jim_thompson5910):

same with 29+25, it should be 29-25

OpenStudy (anonymous):

which makes -4v

OpenStudy (anonymous):

so now I have 2v^2-4v+4

jimthompson5910 (jim_thompson5910):

the stumbling block you're dealing with is taking v^2+10v+25=2v^2+14v+29 and converting it into 0 = v^2 + 4v + 4

OpenStudy (anonymous):

ya

jimthompson5910 (jim_thompson5910):

this is where your mistake occurred

jimthompson5910 (jim_thompson5910):

the key is to get everything to one side

jimthompson5910 (jim_thompson5910):

to do that, we subtract v^2 from both sides, then we subtract 10v from both sides, finally we subtract 25 from both sides

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

why all those subtractions? I did all that to make the left side equal to 0

OpenStudy (anonymous):

so now I have to factor out v^2-4v+4

jimthompson5910 (jim_thompson5910):

more like v^2+4v+4 since 14v-10v = 4v

OpenStudy (anonymous):

I see what I did wrong

jimthompson5910 (jim_thompson5910):

that's great

OpenStudy (anonymous):

would (w-1)^2 be expanded as w^2 +2w+2?

jimthompson5910 (jim_thompson5910):

close, it should be w^2-2w+1

jimthompson5910 (jim_thompson5910):

(w-1)^2 = (w-1)(w-1) = w(w-1) - 1(w-1)

OpenStudy (anonymous):

my online math program logged out and now it changed my problem. I now have to expand (w+6)^2, and expanding that got me x^2+36w+36

jimthompson5910 (jim_thompson5910):

the middle term is off

OpenStudy (anonymous):

I don't know why expanding is so hard for me

jimthompson5910 (jim_thompson5910):

(w+6)^2 = (w+6)(w+6)

OpenStudy (anonymous):

is the middle term 12w?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

I just solved the problem

jimthompson5910 (jim_thompson5910):

that's great, glad you're getting the hang of this

OpenStudy (anonymous):

4